OpenStudy (anonymous):

Integrate 1/ (-v^2-v+3) please

OpenStudy (mathmale):

Looks like a good candidate for the partial fraction method of separating the integrand into easier-to-integrate terms. Are you familiar with partial fractions?

OpenStudy (anonymous):

i don't see how it can be solved using partial fractions. it can't be factored anyways, i wrote down the wrong problem which led me to this mess

OpenStudy (mathmale):

Why not start over again? Post the problem on which you'd like to focus.

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{1}{-v^2-v+3}dv\] right?

OpenStudy (anonymous):

@kimmy0394

OpenStudy (anonymous):

yes sayakdbz

OpenStudy (dumbcow):

hmm how about completing the square, then a trig substitution \[- \int\limits \frac{1}{(v+\frac{1}{2})^2 -\frac{13}{4}}\]

OpenStudy (mathmale):

Kimmy: I'd thought you said you'd posted the wrong problem. Please let those helping you know on what you want to focus.

OpenStudy (anonymous):

ok here's the problem: Use the method discussed under “Homogeneous Equations” to solve (3x^2 -y^2)dx + (xy-x^3y^-1) =0

OpenStudy (anonymous):

a'ight. hold on. u gotta make the denominator a perfect square just as @kimmy0394

OpenStudy (anonymous):

*just as @dumbcow mentioned

OpenStudy (mathmale):

Kimmy: Please warn the others that your focus has shifted to solving this differential equation.

OpenStudy (anonymous):

so i'm just gonna award dumbcow the medal for my original question and i'll close it and post another. Thank you sayakdbz for your input as well